# Bandits with Preference Feedback: A Stackelberg Game Perspective

### v2

@@ 6.1 Benchmarking Confidence Sets
Performance of MaxMinLCB relies on validity and tightness of the LCB. We evaluate the quality of our kernelized confidence bounds, using the potentially simpler task of bandit optimization given logistic feedback. To this end, we fix the acquisition function for the logistic bandit algorithms to the Upper Confidence Bound (UCB) function, and benchmark different methods for calculating the confidence bound. We refer to the algorithm instantiated with the confidence sets of Theorem 2 as LGP-UCB (c.f. Algorithm 2). The Ind-UCB approach assumes that actions are uncorrelated, and maintains an independent confidence interval for each action as in Lattimore and Szepesvári (2020, Algorithm 3). This demonstrates how LGP-UCB utilizes the correlation between actions. We also implement Log-UCB1 (Faury et al., 2020) that assumes that $f$ is a linear function, i.e., $f({\bm{x}})=\theta^{T}{\bm{x}}$ to highlight the improvements gained by kernelization. Last, we compare LGP-UCB with GP-UCB (Srinivas et al., 2010) that estimates probabilities $s(f(\cdot))$ via a kernelized ridge regression task. This comparison highlights the benefits of using our kernelized logistic estimator (Proposition 1) over regression-based approaches (Xu et al., 2020; Kirschner and Krause, 2021; Mehta et al., 2023b, a). Figure 1(a) shows that the cumulative regret of LGP-UCB is the lowest among the baselines. GP-UCB performs closest to LGP-UCB, however, it accumulates regret linearly during the initial steps. Note that GP-UCB and LGP-UCB differ in the estimation of the utility function $f_{t}$ while estimating the width of the confidence bounds similarly. This result suggests that using the logistic-type loss (3) to infer the utility function is advantageous. As expected, Ind-UCB converges at a slower rate than LGP-UCB and GP-UCB due to ignoring the correlation between arms while Log-UCB1’s regret grows linearly as the Ackley function is misspecified under the assumption of linearity. We defer the results on the rest of the utility functions to Table 2 in Appendix E and the figures therein.
@@ 6.2 Benchmarking Acquisition Functions
Hoelder $\boldsymbol{141}\pm 26$ $\boldsymbol{154}\pm 3$ $\boldsymbol{136}\pm 15$ $204\pm 20$ $200\pm 28$ $\boldsymbol{132}\pm 15$
@@
In this section, we compare MaxMinLCB with other utility-based bandit algorithms. To isolate the benefits of our acquisition function, we instantiate all algorithms with the same confidence sets, and use our improved preferred-based bound of Corollary 5. Therefore, our implementation differs from the corresponding references, while we refer to the algorithms by their original name. We consider the following baselines. Doubler and MultiSBM (Ailon et al., 2014) choose ${\bm{x}}_{t}$ as a reference action from the recent history of actions and pair it with ${\bm{x}}^{\prime}_{t}$ which maximizes the joint UCB (cf. Algorithm 5 and 6). RUCB (Zoghi et al., 2014a) chooses ${\bm{x}}^{\prime}_{t}$ similarly, however, it selects the reference action uniformly at random from ${\mathcal{M}}_{t}$ (Algorithm 7). MaxInP (Saha, 2021) also maintains the set of plausible maximizers ${\mathcal{M}}_{t}$, and at each time step, it selects the pair of actions that maximize $\sigma_{t}^{D}({\bm{x}},{\bm{x}}^{\prime})$ (Algorithm 3). IDS (Kirschner and Krause, 2021) selects the reference action greedily by maximizing $f_{t}$, and pairs it with an informative action (Algorithm 4). Notably, all algorithms, with the exception of MaxInP, choose one of the actions independently and use it as a reference point when selecting the other one. Figure 3 illustrates the differences in action selection between UCB, maximum information, and MaxMinLCB approaches. We note that POP-BO (Xu et al., 2024) and MultiSBM only differ in the estimation of the confidence set. Since we deploy the same confidence set for all acquisition functions, the two algorithms are equivalent and we use MultiSBM in our results, however, comparisons hold for POP-BO as well.
@@
Figure 1(b) benchmarks the algorithms using the Ackley utility function, where MaxMinLCB outperforms the baselines. All algorithms suffer from close-to-linear regret during the initial stages of learning, suggesting that there is an inevitable exploration phase. Notably, MaxMinLCB, IDS, and Doubler are the first to select actions with high utility, while RUCB and MaxInP explore for longer. Table 1 shows the dueling regret for all utility functions. MaxMinLCB consistently outperforms the baselines across the analyzed functions and achieves a low standard error, supporting its efficiency in balancing exploration and exploitation in the preference feedback setting. While MultiSBM is consistently outperformed, we do not observe a clear ranking among the rest of the baselines. For instance, IDS achieves the smallest regret for optimizing Matyas, while RUCB excels on the Branin function. This indicates the challenges each function offers and the performance of the action selection is task dependent. The consistent performance of MaxMinLCB demonstrates its robustness against the underlying unknown utility function.
@@ 6.3 Real-world Experiment
To further demonstrate the scalability and applicability of MaxMinLCB, we conduct an experiment on the Yelp open dataset of restaurant reviews that we use to design a complex real-world problem. The goal of this experiment is to identify restaurants for each user that fit their respective preferences the best via sequential recommendations. The final dataset after cleaning consists of $275$ restaurants and $20$ users. We define the action space ${\mathcal{X}}$ by assigning to each restaurant their respective $32$-dimensional embedding of their reviews, i.e., ${\mathcal{X}}\subseteq{\mathbb{R}}^{32}$. The dataset provides utility values for users in the form of ratings on the scale of $1$ to $5$, however, not all users rated every restaurant. We estimate missing ratings using collaborative filtering (Schafer et al., 2007). Further details on the data processing are deferred to Section D.1.
Figure 2 shows that the results of this larger problem align with previous conclusions. MaxMinLCB remains the best-performing algorithm with MultiSBM following second. The cumulative regret is also reported in Table 1. Note that neither of the algorithms is tuned or modified for this experiment. These results are only intended to demonstrate that 1) the computations easily scale, and 2) the kernelized approach is still applicable in a text-based domain, by using high-quality vector embeddings.
Figure 2: Regret of learning restaurant recommendations using the Yelp open dataset with preference feedback. Algorithms with different acquisition functions, all using our confidence sets. MaxMinLCB is more sample-efficient.
@@ 7 Conclusion
We thank Scott Sussex for the thorough feedback on the manuscript. This research was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and Innovation Program Grant agreement No. 815943. Barna Pásztor was supported by an ETH AI Center doctoral fellowship.
@@ Appendix A Proofs for Bandits with Logistic Feedback
We have written the equations in the main text in terms of kernel matrices and function evaluations, for easier readability. For the purpose of the proof however, we mainly rely on entities in the Hilbert space. Consider the operator ${\bm{\phi}}:{\mathcal{X}}\rightarrow{\mathcal{H}}$ which corresponds to kernel $k$ and satisfies $k({\bm{x}},\cdot)={\bm{\phi}}({\bm{x}})$. Then by Mercer’s theorem, any $f\in{\mathcal{H}}_{k}$ may be written as $f={\bm{\theta}}^{\top}{\bm{\phi}}$, where ${\bm{\theta}}\in\ell_{2}({\mathbb{N}})$ and has a $B$-bounded $\ell_{2}$ norm. For a sequence of points ${\bm{x}}_{1},\dots,{\bm{x}}_{t}\in{\mathcal{X}}$, we define the (infinite-dimensional) feature map $\Phi_{t}=\left[{\bm{\phi}}({\bm{x}}_{1}),\cdots,{\bm{\phi}}({\bm{x}}_{t})\right]^{\top}$, which gives rise to the kernel matrix $K_{t}:{\mathbb{R}}^{t}\rightarrow{\mathbb{R}}^{t}$ and the covariance operator $S_{t}:{\mathcal{H}}_{k}\rightarrow{\mathcal{H}}_{k}$, respectively defined as $K_{t}=\Phi_{t}\Phi_{t}^{\top}$ and $S_{t}=\Phi_{t}^{\top}\Phi_{t}$. Let ${\bm{I}}_{t}$ denote the $t$-dimensional identity matrix, and ${\bm{I}}_{\mathcal{H}}$ be the identity operator acting on the RKHS. Then it is widely known that the covariance and kernel operators are connected via $\det({\bm{I}}_{\mathcal{H}}+\rho^{-2}S_{t})=\det({\bm{I}}_{t}+\rho^{-2}K_{t})$ for any $t\geq 1$ and $\rho\neq 0$. For operators on the Hilbert space, $\det(A)$ refer to a Fredholm determinant [c.f. Lax, 2002]. Lastly, in the appendix, we refer to the true unknown utility function as $f^{\star}({\bm{x}})={\bm{\phi}}^{\top}({\bm{x}}){\bm{\theta}}^{\star}$. In the main text, the true utility is simply referred to as $f$.
@@
Our analysis relies on a concentration bound for ${\mathcal{H}}$-valued martingale sequences stated in Abbasi-Yadkori [2013] and later in Whitehouse et al. [2023]. Below, we have adapted the statement to match our notation.
Lemma 7 (Corollary 1 Whitehouse et al. [2023]).
@@
The following lemma, which extends Faury et al. [2020, Lemma 8] to ${\mathcal{H}}$-valued operators, expresses the closeness of ${\bm{\theta}}_{t}$ and ${\bm{\theta}}^{\star}$ in the gradient space, with respect to the norm of the covariance matrix.
@@
Set $10<\delta<$ 1. Under the assumptions of Theorem 2
@@
where $\varepsilon_{\tau}=y_{\tau}-s({\bm{\phi}}^{\top}({\bm{x}}_{\tau}){\bm{\theta}}^{\star})$ is a history dependent random variable in $[0,1]$ due to the data model. To bound the first term, we recognize that any random variable in $[0,1]$ is $\sigma$ sub-Gaussian with $\sigma=0.5$ and apply Lemma 7. We obtain that for all $t\geq 0$, with probability greater than $1-\delta$
@@
Set $10<\delta<$ 1 and consider the confidence sets
@@
Lastly, we prove a formal extension of Theorem 2.
Theorem 10 (Theorem 2 - Formal).
Set $10<\delta<$ 1 and consider the confidence sets ${\mathcal{E}}_{t}(\delta)\subset{\mathcal{H}}_{k}$
@@
Initialize Set $(\beta_{t})_{t\geq 1}$ according to Theorem 2.
@@
     Calculate ${f_{t}}$ acc. to Proposition 1 and update $\sigma_{t}$ acc. to Theorem 2.
@@
Given the confidence set of Theorem 2, we give extend the LGP-UCB algorithm of Faury et al. to the kernelized setting (c.f. Algorithm 2) and prove that it satisfies sublinear regret.
Proof of Corollary 3.
@@
Then using Theorem 10, we obtain the following for the regret at step $t$,
@@
Appendix B Helper Lemmas for Appendix A
@@
$\left\lVert{\bm{\theta}}_{1}-{\bm{\theta}}_{2}\right\rVert_{V_{t}}$ (11)${{\leq}}\sqrt{\kappa}\left\lVert{\bm{\theta}}_{1}-{\bm{\theta}}_{2}\right\rVert_{G_{t}({\bm{\theta}}_{1},{\bm{\theta}}_{2})}$
(12)${{\leq}}\sqrt{\kappa}\left\lVert g_{t}({\bm{\theta}}_{1})-g_{t}({\bm{\theta}}_{2})\right\rVert_{G_{t}^{-1}({\bm{\theta}}_{1},{\bm{\theta}}_{2})}$
(11)${{\leq}}\kappa\left\lVert g_{t}({\bm{\theta}}_{1})-g_{t}({\bm{\theta}}_{2})\right\rVert_{V_{t}^{-1}}$
@@
The following two lemmas are standard results in kernelized bandits [Srinivas et al., 2010, Chowdhury and Gopalan, 2017, e.g.,]. We include it here for completeness.
@@
=(13)${{=}}\Phi_{t}^{\top}\left(\Phi_{t}\Phi_{t}^{\top}+\lambda\kappa{\bm{I}}_{t}\right)^{-1}{\bm{k}}_{t}({\bm{x}})+\lambda\kappa\left(\Phi_{t}^{\top}\Phi_{t}+\lambda\kappa{\bm{I}}\right)^{-1}{\bm{\phi}}({\bm{x}}).$ ) .
@@
=(14)${{=}}{\bm{k}}_{t}^{\top}({\bm{x}})\left(K_{t}+\lambda\kappa{\bm{I}}_{t}\right)^{-1}{\bm{k}}_{t}({\bm{x}})+\lambda\kappa{\bm{\phi}}^{\top}({\bm{x}})V_{t}^{-1}{\bm{\phi}}({\bm{x}})$ )
To obtain the third equation we have used the fact that for bounded operators on Hilbert spaces, the inverse of the adjoint is equal to the adjoint of the inverse [e.g., Theorem 10.19, Axler, 2020]. Re-ordering the equation above we obtain $\sigma_{t}^{2}({\bm{x}})=\lambda\kappa\left\lVert{\bm{\phi}}({\bm{x}})\right\rVert_{V_{t}^{-1}}^{2}$, concluding the proof. ∎
@@
Proof of Lemma 14.
By Srinivas et al. [2010, Lemma 5.3],
@@
Following the technique in Srinivas et al. [2010, Lemma 5.4], since $\sigma^{2}_{t}\leq 1$, then $(\lambda\kappa)^{-1}\sigma^{2}_{t}\in[0,(\lambda\kappa)^{-1}]$. Now for any $z\in[0,(\lambda\kappa)^{-1}]$, $z\leq C\log(1+z)$ where $C=1/(\lambda\kappa\log(1+(\lambda\kappa)^{-1}))$. We then may write,

### v3

@@ 6.1 Benchmarking Confidence Sets
Performance of MaxMinLCB relies on validity and tightness of the LCB. We evaluate the quality of our kernelized confidence bounds, using the potentially simpler task of bandit optimization given logistic feedback. To this end, we fix the acquisition function for the logistic bandit algorithms to the Upper Confidence Bound (UCB) function, and benchmark different methods for calculating the confidence bound. We refer to the algorithm instantiated with the confidence sets of Theorem˜2 as LGP-UCB (c.f. Algorithm˜2). The Ind-UCB approach assumes that actions are uncorrelated, and maintains an independent confidence interval for each action as in lattimore2020bandit. This demonstrates how LGP-UCB utilizes the correlation between actions. We also implement Log-UCB1 (faury2020improved) that assumes that $f$ is a linear function, i.e., $f({\bm{x}})=\theta^{T}{\bm{x}}$ to highlight the improvements gained by kernelization. Last, we compare LGP-UCB with GP-UCB (srinivas2009gaussian) that estimates probabilities $s(f(\cdot))$ via a kernelized ridge regression task. This comparison highlights the benefits of using our kernelized logistic estimator (Proposition˜1) over regression-based approaches (xu2020zeroth; kirschner2021bias; mehta2023kernelized; mehta2023sample). Figure˜1(a) shows that the cumulative regret of LGP-UCB is the lowest among the baselines. GP-UCB performs closest to LGP-UCB, however, it accumulates regret linearly during the initial steps. Note that GP-UCB and LGP-UCB differ in the estimation of the utility function $f_{t}$ while estimating the width of the confidence bounds similarly. This result suggests that using the logistic-type loss (3) to infer the utility function is advantageous. As expected, Ind-UCB converges at a slower rate than LGP-UCB and GP-UCB due to ignoring the correlation between arms while Log-UCB1’s regret grows linearly as the Ackley function is misspecified under the assumption of linearity. We defer the results on the rest of the utility functions to Table˜2 in Appendix˜E and the figures therein.
@@ 6.2 Benchmarking Acquisition Functions
Hoelder $\boldsymbol{141}\pm 26$ $\boldsymbol{154}\pm 3$ $\boldsymbol{136}\pm 15$ $4\pm 20 200±2$$0\pm 28$ $\boldsymbol{132}\pm 15$
@@
In this section, we compare MaxMinLCB with other utility-based bandit algorithms. To isolate the benefits of our acquisition function, we instantiate all algorithms with the same confidence sets, and use our improved preferred-based bound of Corollary˜5. Therefore, our implementation differs from the corresponding references, while we refer to the algorithms by their original name. We consider the following baselines. Doubler and MultiSBM (ailon2014reducing) choose ${\bm{x}}_{t}$ as a reference action from the recent history of actions and pair it with ${\bm{x}}^{\prime}_{t}$ which maximizes the joint UCB (cf. Algorithm˜5 and 6). RUCB (zoghi2014relative) chooses ${\bm{x}}^{\prime}_{t}$ similarly, however, it selects the reference action uniformly at random from ${\mathcal{M}}_{t}$ (Algorithm˜7). MaxInP (saha2021optimal) also maintains the set of plausible maximizers ${\mathcal{M}}_{t}$, and at each time step, it selects the pair of actions that maximize $\sigma_{t}^{D}({\bm{x}},{\bm{x}}^{\prime})$ (Algorithm˜3). IDS (kirschner2021bias) selects the reference action greedily by maximizing $f_{t}$, and pairs it with an informative action (Algorithm˜4). Notably, all algorithms, with the exception of MaxInP, choose one of the actions independently and use it as a reference point when selecting the other one. Figure˜3 illustrates the differences in action selection between UCB, maximum information, and MaxMinLCB approaches. We note that POP-BO (xu2024principled) and MultiSBM only differ in the estimation of the confidence set. Since we deploy the same confidence set for all acquisition functions, the two algorithms are equivalent and we use MultiSBM in our results, however, comparisons hold for POP-BO as well.
@@
Figure˜1(b) benchmarks the algorithms using the Ackley utility function, where MaxMinLCB outperforms the baselines. All algorithms suffer from close-to-linear regret during the initial stages of learning, suggesting that there is an inevitable exploration phase. Notably, MaxMinLCB, IDS, and Doubler are the first to select actions with high utility, while RUCB and MaxInP explore for longer. Table˜1 shows the dueling regret for all utility functions. MaxMinLCB consistently outperforms the baselines across the analyzed functions and achieves a low standard error, supporting its efficiency in balancing exploration and exploitation in the preference feedback setting. Only MultiSBM shows comparable performance to MaxMinLCB and even outperforms it on the Matyas function which is a relatively smooth function posing a simple optimization problem. However, MaxMinLCB achieves lower regret on the Ackley and Eggholder functions which obtain many local optima and steeper gradients showing that MaxMinLCB is preferable for challenging optimization problems. Other acquisition functions work well only in certain cases, e.g., IDS achieves the smallest regret for optimizing Matyas, while RUCB excels on the Branin function. This indicates the challenges each utility function offers and the performance of the action selection is task dependent. The consistent performance of MaxMinLCB demonstrates its robustness against the underlying unknown utility function.
@@ 6.3 Real-world Experiment
To demonstrate the scalability and applicability of MaxMinLCB, we conduct an experiment on the Yelp dataset of restaurant reviews, which consists of $275$ restaurants and $20$ users after the pre-processing stage. For each user, we want to find the restaurants that best fit their preference, via making sequential recommendations and receiving comparative feedback. We define the action space $\mathcal{X}$} by assigning to each restaurant their respective $32$-dimensional embedding of their reviews, i.e., ${\mathcal{X}}\subseteq{\mathbb{R}}^{32}$. The dataset provides utility values for users in the form of ratings on the scale of $1$ to $5$, however, not all users rated every restaurant.
We estimate missing ratings using collaborative filtering (schafer2007collaborative). Further details on the data processing are deferred to Section˜D.1. Figure˜2 shows that the results of this larger problem align with previous conclusions. MaxMinLCB remains the best-performing algorithm with MultiSBM following second. The cumulative regret is also reported in Table˜1. Note that neither of the algorithms is tuned or modified for this experiment. These results are only intended to demonstrate that 1) the computations easily scale, and 2) the kernelized approach is still applicable in a text-based domain, by using high-quality vector embeddings.
Figure 2: LGP-UCB is more sample-efficient when making restaurant recommendations based on Yelp open dataset with preference feedback. All baselines use the confidence sets of Cor. 5.
@@ 7 Conclusion
We thank Scott Sussex for his thorough feedback. This research was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and Innovation Program Grant agreement No. 815943. Barna Pásztor was supported by an ETH AI Center doctoral fellowship, and Parnian Kassraie by a Google PhD Fellowship.
@@ Appendix A Proofs for Bandits with Logistic Feedback
We have written the equations in the main text in terms of kernel matrices and function evaluations, for easier readability. For the purpose of the proof however, we mainly rely on entities in the Hilbert space. Consider the operator ${\bm{\phi}}:{\mathcal{X}}\rightarrow{\mathcal{H}}$ which corresponds to kernel $k$ and satisfies $k({\bm{x}},\cdot)={\bm{\phi}}({\bm{x}})$. Then by Mercer’s theorem, any $f\in{\mathcal{H}}_{k}$ may be written as $f={\bm{\theta}}^{\top}{\bm{\phi}}$, where ${\bm{\theta}}\in\ell_{2}({\mathbb{N}})$ and has a $B$-bounded $\ell_{2}$ norm. For a sequence of points ${\bm{x}}_{1},\dots,{\bm{x}}_{t}\in{\mathcal{X}}$, we define the (infinite-dimensional) feature map $\Phi_{t}=\left[{\bm{\phi}}({\bm{x}}_{1}),\cdots,{\bm{\phi}}({\bm{x}}_{t})\right]^{\top}$, which gives rise to the kernel matrix $K_{t}:{\mathbb{R}}^{t}\rightarrow{\mathbb{R}}^{t}$ and the covariance operator $S_{t}:{\mathcal{H}}_{k}\rightarrow{\mathcal{H}}_{k}$, respectively defined as $K_{t}=\Phi_{t}\Phi_{t}^{\top}$ and $S_{t}=\Phi_{t}^{\top}\Phi_{t}$. Let ${\bm{I}}_{t}$ denote the $t$-dimensional identity matrix, and $\bm{I}$$_{\mathcal{H}}$ be the identity operator acting on the RKHS. Then it is widely known that the covariance and kernel operators are connected via $\det({\bm{I}}_{\mathcal{H}}+\rho^{-2}S_{t})=\det({\bm{I}}_{t}+\rho^{-2}K_{t})$ for any $t\geq 1$ and $\rho\neq 0$. For operators on the Hilbert space, $\det(A)$ refer to a Fredholm determinant [c.f. lax2002functional]. Lastly, in the appendix, we refer to the true unknown utility function as $f^{\star}({\bm{x}})={\bm{\phi}}^{\top}({\bm{x}}){\bm{\theta}}^{\star}$. In the main text, the true utility is simply referred to as $f$.
@@
Our analysis relies on a concentration bound for $\mathcal{H}$}-valued martingale sequences stated in abbasi2013online and later in whitehouse2023improved. Below, we have adapted the statement to match our notation.
Lemma 7 (Corollary 1 whitehouse2023improved).
@@
The following lemma, which extends faury2020improved to $\mathcal{H}$}-valued operators, expresses the closeness of ${\bm{\theta}}_{t}$ and $\bm{\theta}$$^{\star}$ in the gradient space, with respect to the norm of the covariance matrix.
@@
Set $0<\delta<1$. Under the assumptions of Theorem˜2
@@
where $\varepsilon_{\tau}=y_{\tau}-s({\bm{\phi}}^{\top}({\bm{x}}_{\tau}){\bm{\theta}}^{\star})$ is a history dependent random variable in $[0,1]$ due to the data model. To bound the first term, we recognize that any random variable in $[0,1]$ is $\sigma$ sub-Gaussian with $\sigma=0.5$ and apply Lemma˜7. We obtain that for all $t\geq 0$, with probability greater than $1-\delta$
@@
Set $0<\delta<1$ and consider the confidence sets
@@
Lastly, we prove a formal extension of Theorem˜2.
Theorem 10 (Theorem˜2 - Formal).
Set $0<\delta<1$ and consider the confidence sets ${\mathcal{E}}_{t}(\delta)\subset{\mathcal{H}}_{k}$
@@
Initialize Set $(\beta_{t})_{t\geq 1}$ according to Theorem˜2.
@@
  Calculate $f_{t}$} acc. to Proposition˜1 and update $\sigma_{t}$ acc. to Theorem˜2.
@@
Given the confidence set of Theorem˜2, we give extend the LGP-UCB algorithm of faury2020improved to the kernelized setting (c.f. Algorithm˜2) and prove that it satisfies sublinear regret.
Proof of Corollary˜3.
@@
Then using Theorem˜10, we obtain the following for the regret at step $t$,
@@
Appendix B Helper Lemmas for Appendix˜A
@@
[…]
@@
The following two lemmas are standard results in kernelized bandits [srinivas2009gaussian, chowdhury2017kernelized, e.g.,]. We include it here for completeness.
@@
[…]
@@
[…]
To obtain the third equation we have used the fact that for bounded operators on Hilbert spaces, the inverse of the adjoint is equal to the adjoint of the inverse [e.g., Theorem 10.19, axler2020measure]. Re-ordering the equation above we obtain $\sigma_{t}^{2}({\bm{x}})=\lambda\kappa\left\lVert{\bm{\phi}}({\bm{x}})\right\rVert_{V_{t}^{-1}}^{2}$, concluding the proof. ∎
@@
Proof of Lemma˜14.
By srinivas2009gaussian,
@@
Following the technique in srinivas2009gaussian, since $\sigma^{2}_{t}\leq 1$, then $(\lambda\kappa)^{-1}\sigma^{2}_{t}\in[0,(\lambda\kappa)^{-1}]$. Now for any $z\in[0,(\lambda\kappa)^{-1}]$, $z\leq C\log(1+z)$ where $C=1/(\lambda\kappa\log(1+(\lambda\kappa)^{-1}))$. We then may write,